OpenStudy (anonymous):

One diagonal of a cyclic quadrilateral coincides with a diameter of a circle whose area is 36 pi cm^2. If the other diagonal which measures 8 cm meets the first diagonal at the right angles,find the area of the quadrilateral. Please help me :(

OpenStudy (perl):



OpenStudy (anonymous):

what's next how can i find its area? :( please help me

OpenStudy (anonymous):

master @perl what's next? :)

OpenStudy (anonymous):

@perl

OpenStudy (perl):

you to MIT ?

OpenStudy (anonymous):

mapua :)

OpenStudy (perl):

whats mapua?

OpenStudy (anonymous):

i mean can you guve me the solution for the problem @@perl ? :( =))

OpenStudy (perl):

sure, one sec

OpenStudy (perl):

If a quadrilateral has perpendicular diagonals, d1 and d2, then area=(d1 x d2)/2

OpenStudy (anonymous):

thank youu in advance!!! you are an angel! haha =)))

OpenStudy (perl):

Pi * r^2 = 36*Pi r^2 = 36 r = 6 now diameter = 2*r so diameter = 2*6 = 12

OpenStudy (perl):

now use this theorem: if a quadrilateral has perpendicular diagonals, d1 and d2, then area=(d1 x d2)/2 Area = ( 12 x 8 ) / 2 = 48

OpenStudy (perl):

so the answer is 48

OpenStudy (perl):

please check and tell me if that is correct

OpenStudy (anonymous):

is that the actual image or illustration for the problem? @perl

OpenStudy (perl):

it is an example

OpenStudy (anonymous):

where did you get the d1x d2/2 ? master @perl ? :))